Solve this equation by using the quadratic form.
x(1/2) - 5 sqrt 3x(1/4) + 18 = 0

Answer by stanbon(75887)

Solve this equation by using the quadratic form.
x(1/2) - 5 sqrt 3x(1/4) + 18 = 0
-----
Let w = x^(1/4)
Then w^2 = x^(1/2)
-------
Comment:
I think you have not posted your problem correctly.
==========
Cheers,
Stan H.